OpenStudy (fanduekisses):

pls help. ;'( use the properties of logarithms to rewrite and simplify the logarithmic expression.

OpenStudy (fanduekisses):

\[\log _{5}(\frac{ 1 }{ 250 })\]

OpenStudy (fanduekisses):

\[\log_{5}(1)-\log_{5}(250)\]???

OpenStudy (fanduekisses):

@amistre64 @dan815 @satellite73

OpenStudy (fanduekisses):

then what?

OpenStudy (amistre64):

the question seems vague. what constitutes a 'simplified' version of it?

OpenStudy (amistre64):

what does log(1) always equal?

OpenStudy (amistre64):

so 250 a multiple of 5 by chance?

OpenStudy (amistre64):

what log properties do you know?

OpenStudy (fanduekisses):

yes , 50 right

OpenStudy (fanduekisses):

ohh could I use the change of base ? lol

OpenStudy (mertsj):

log(base5) of1 = 0 so the simplified version is - log(base5)250

OpenStudy (fanduekisses):

The answer key says it's -3-log(base 5) (2) but how?

OpenStudy (amistre64):

notifs didnt say nothing about 15 minutes ago ...

OpenStudy (amistre64):

what are your log properties?

OpenStudy (fanduekisses):

Power, product, and quotient

OpenStudy (mertsj):

Or perhaps this: \[\log_{5}1-\log_{5}5^3(2)=0-(\log_{5}5^3 +\log_{5}2)=-3-\log_{5}2 \]

OpenStudy (fanduekisses):

omg, that makes so much sense, I see what was done there. :D

OpenStudy (fanduekisses):

Pretty much all the log properties were used there haha. :)